julie and kristen are the partners in a local sporting good shop . they needed $51,000 to start the buisness. they invested in the ratio 5:12 respectively. what percent of the business was owned by Kristen ? if the business grows to $3,000,000 what percent of it will julie own ?
Answer by stanbon(75887)

julie and kristen are the partners in a local sporting good shop . they needed $51,000 to start the buisness. they invested in the ratio 5:12 respectively. what percent of the business was owned by Kristen ? if the business grows to $3,000,000 what percent of it will julie own ?
----
5:12 is the same as 5x:12x where Kristen's share is 12x
---
Julie's percentage = 5x/17x = 5/17 = 29.4%
---
Kristen's percentage = 12x/17x = 12/17 = 70.6%
==========================
Cheers,
Stan H.
==========================